1. The odds that a person A speaks the truth is 8:9 and the odds that person B speaks the truth is 5:7. In what percentage of cases are they likely to contradict each other on an identical point 
1
Expert's answer
2020-07-09T19:57:15-0400

P(A speaks the truth)= "\frac{8}{8+9}= \frac{8}{17}"

P(A speaks a lie)= "\frac{9}{8+9}= \frac{9}{17}"

P(B speaks the truth)= "\frac{5}{5+7}= \frac{5}{12}"

P(B speaks a lie)="\frac{7}{5+7}= \frac{7}{12}"

They will contradict each other when A tells the truth/ B tells a lie or A tells a lie /B tells the truth


P(contradict)="\frac{8}{17}*\frac{7}{12} +\frac{9}{17}*\frac{5}{12}"


="\frac{56}{204}+\frac{45}{204}"


="\frac{101}{204}"

=0.495.

hence the percentage is 49.5%.